OpenStudy (anonymous):

Find the slope of the line passing through the points (-3, -10) and (4, 5)

OpenStudy (anonymous):

help me solve please?

OpenStudy (anonymous):

the equation for the slope looks like (y1 - y2) m = -------- (x1 - x2) so if you wanted to find the slope between the points (-2,3) and (1,4) then you would do (3 - 4) m = -------- = 1/3 (-2 - 1) just apply this to your problem

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

glad I could help

OpenStudy (anonymous):

You helped a lot!